OpenStudy (anonymous):

4+2/x / x/3+1/6 can someone help me solve this please

OpenStudy (amistre64):

\[\frac{\frac{4+2}{x}}{\frac{x}{3}+\frac{1}{6}}\]this thing?

OpenStudy (anonymous):

is there and "=" sign or is this a compound fraction?

OpenStudy (amistre64):

solve might mean simplify ;) just my assumption tho

OpenStudy (anonymous):

\[4 + 2/x \div x/3 + 1/6\]

OpenStudy (anonymous):

in that case multiply top and bottom by 6x to clear fractions. or if you want you can do the subtraction in the denominator, and then divide by multiplying by the reciprocal of the denominator (after you subtract).

OpenStudy (amistre64):

multiply everything by 6x could work

OpenStudy (anonymous):

ooooh \[4+\frac{2}{x} \div \frac{x}{3} + \frac{1}{6}\] ?

OpenStudy (anonymous):

yes that

OpenStudy (mathmagician):

12/x

OpenStudy (anonymous):

can you tell me how you got the answer please

OpenStudy (anonymous):

\[4+\frac{2}{x} \times \frac{3}{x} + \frac{1}{6}=4+ \frac{1}{6}+\frac{6}{x^2}=\frac{25}{6}+\frac{6}{x^2}\]

OpenStudy (amistre64):

24x+12 12(2x+1) ------- = -------- = 4 if i did it right :) 6x +3 3(2x+1)

OpenStudy (mathmagician):

\[4+2/x=(4x+2)/x and x/3+1/6=(2x+1)/6. Then 4x+2/6 multiplied by 6/(2x+1) is equal \to 12/x\]

OpenStudy (anonymous):

sorry i did not realize it was \[\frac{\frac{x+2}{x}}{\frac{x}{3}+\frac{1}{6}}\]

OpenStudy (amistre64):

but i see an error in mine :) 6x(x/3) = 2x^2

OpenStudy (amistre64):

24x +12 -------- perhaps? 2x^2 +x

OpenStudy (anonymous):

it it perhaps \[\frac{4+\frac{2}{x}}{\frac{x}{3}+\frac{1}{6}}\]?

OpenStudy (amistre64):

\[\frac{4x+2}{x}*\frac{6}{2x+1}\]

OpenStudy (amistre64):

12/x